# A Comprehensive Survey on Multi-modal Conversational Emotion Recognition with Deep Learning

### v1

@@ 7.2. Public Opinion Analysis
Multi-modal conversational emotion analysis also has a wide application value in the field of opinion mining, which can help mine and analyze people’s opinions and emotions expressed in text, video and audio (Tan et al., 2021). For example, in market research, researchers can use emotion analysis techniques to analyze users’ opinions on different products, as well as users’ purchase intentions. This has great potential value for developing marketing strategies.
@@ 7.3. Recommendations Systems
Multi-modal conversational emotion analysis in recommender systems can help personalize recommendations more in line with users’ emotions and preferences (Ding et al., 2022). For example, the recommendation system can recommend products that users are more interested in according to the emotional changes of consumers, and can perform emotion analysis on multi-modal data of user evaluations to realize real-time early warning and disposal of negative product evaluations.
@@ 7.4. Medical Care
Multi-modal conversational emotion analysis plays an important role in many aspects in the field of health care (Saganowski et al., 2022). It can help medical institutions and doctors better understand the current emotional state of patients, so as to give better treatment plans. For example, doctors can use emotion analysis technology to analyze patients’ medical records and symptom descriptions, so as to better understand the patient’s emotional state and help make more accurate diagnosis and treatment plans. Furthermore, in diagnosis and treatment decisions, understanding a patient’s emotional state is important for developing an appropriate medical regimen. Emotion analysis can help doctors make more targeted decisions.
@@ 7.5. Financial Field Analysis
In the field of financial analysis, emotion analysis can help financial practitioners and investors better understand the emotional state of the market and predict market trends, thereby helping investors make correct investment decisions (Gerczuk et al., 2021). For example, some financial institutions and financial practitioners use emotion analysis technology to calculate market emotion index to measure market emotion, which can provide investors with an objective reference value.
@@ 7.6. Social Robot
Multimodal conversational emotion recognition has many potential applications on social robots, which can enhance the capabilities of social robots and make them more intelligent and humane (Lee et al., 2022). Social robots can use multimodal emotion recognition to sense the emotional state of the users they interact with. This includes identifying users’ facial expressions, voice emotions, text emotions, and other modal emotional signals (Laban et al., 2022). The robot can then adjust its interaction to better meet the user’s emotional needs, providing support, comfort or entertainment. In addition, social robots can use MCER to better understand users’ needs and emotional states to provide personalized suggestions and assistance. For example, in the field of mental health, robots can provide targeted psychological support suggestions based on the user’s emotional state.
@@ 8.2. Data is Heterogeneous and Noisy
Multi-modal conversation emotion recognition models need to fully eliminate heterogeneity and noise information between modalities to achieve accurate prediction or classification of emotions. Multi-modal data is naturally heterogeneous, and features of different modalities have huge differences in processing methods and representation forms. In addition, multi-modal conversation data contains a large amount of redundant or noisy information, and its emotion is usually determined by only a small amount of consistent key information, such as certain words in a sentence, a specific frequency band in speech, or a particular expression in a video. Even in some extreme cases, part of the modal information is basically unavailable under noise interference, such as ambiguous sentence expressions, noise in the speech, blocked expressions, etc. Therefore, the heterogeneity and noise of data limit the effectiveness of current multi-modal conversational emotion recognition models.
@@ 8.4. Consistent Semantic Association
Multi-modal conversation emotion recognition requires the model to fully learn the consistent semantics between modalities to filter noise information and eliminate heterogeneity between modalities, which is the basis for building an accurate multimodal dialogue emotion recognition model. However, the consistent semantic association of multi-modal conversation is more complicated and is not only related to the multi-modal context but also to factors such as the conversation scene, the speaker’s own emotional inertia, and the speaker’s stimulation. In addition, multi-modal data are heterogeneous, each modality has differentiated representation and distribution characteristics in space, and some consistent semantic associations are hidden in the feature distribution space between modalities. Therefore, efficiently performing consistent semantic association is the primary issue that needs to be considered at the model level.
@@ 11. Acknowledgments
This work is supported by National Natural Science Foundation of China (Grant No. 62372478, No. 61802444); the Research Foundation of Education Bureau of Hunan Province of China (Grant No. 22B0275, No.20B625).

### v2

@@ 7.2. Public Opinion Analysis
Multi-modal conversational emotion analysis also has a wide application value in opinion mining, which can help mine and analyze people’s opinions and emotions expressed in text, video and audio (Tan et al., 2021). For example, emotion analysis of online public opinion on emergencies can better understand public emotion and conduct effective crisis management. ECR-BERT (Wan et al., 2024) proposed a BERT-based model that integrates emotion-cognitive reasoning mechanisms, enabling more accurate emotional understanding in complex scenarios such as public opinion analysis during sudden events. Compared with the standard BERT model, ECR-BERT achieved absolute average accuracy improvements of 0.82%, 1.74%, 0.98%, and 1.37% across different datasets. This enhanced emotional recognition capability helps to more precisely capture public sentiment dynamics, providing valuable support for timely public opinion monitoring and effective crisis management.
@@ 7.3. Recommendations Systems
Multi-modal conversational emotion analysis in recommender systems can help personalize recommendations more in line with users’ emotions and preferences (Ding et al., 2022). For example, the recommendation system can recommend products that users are more interested in according to the emotional changes of consumers, and can perform emotion analysis on multi-modal data of user evaluations to realize real-time early warning and disposal of negative product evaluations. For example, Agent4Rec (Zhang et al., 2024b) displays four movies on each recommendation page, and the agent will decide whether to continue to the next recommendation page or exit the recommendation system based on their satisfaction. After the agent exits, the system will ask him to give a satisfaction score of the recommendation system, ranging from 1 to 10. Ratings above 3 are regarded as signals of liking. After the entire simulation is completed, the following multi-faceted indicators are collected: average viewing rate, average number of likes, average like ratio, average number of exit pages, and average user satisfaction score. The satisfaction score of random recommendations is 2.93, while the satisfaction score of the recommendation algorithm based on emotional preferences is 3.85. The experimental results prove the effectiveness of emotion recognition technology in the field of social media.
@@ 7.4. Medical Care
Multi-modal conversational emotion analysis plays an important role in many aspects in the field of health care (Saganowski et al., 2022). It can help medical institutions and doctors better understand the current emotional state of patients, so as to give better treatment plans. For example, by performing emotion analysis on unstructured data such as patients’ medical records, consultation conversations, online consultation texts, or social media posts, the medical system can identify patients’ potential emotional states, thereby providing supplementary psychological information for clinical diagnosis (Sun et al., 2024). In addition, in mental health assessments, emotion analysis can be used to detect early signs of depression tendencies or anxiety symptoms, which helps to achieve early screening and early intervention. In chronic disease management, the system can continuously monitor patients’ emotional responses to treatment plans and help doctors dynamically adjust treatment strategies.
@@ 7.5. Financial Field Analysis
In the field of financial analysis, emotion analysis can help financial practitioners and investors better understand the emotional state of the market and predict market trends, thereby helping investors make correct investment decisions (Gerczuk et al., 2021). For example, the improved GPT model (Xing, [n.d.]) combined with the emotion analysis module achieved an accuracy of 88.34% in the financial emotion classification task, which is significantly better than the original GPT model (74.57% accuracy) that did not use emotion information. The results show that emotion factors can provide the model with richer semantics and market tendency judgment basis, further verifying the actual effectiveness and research value of emotion analysis in financial text understanding and trend prediction.
@@ 7.6. Social Robot
Multimodal conversational emotion recognition has many potential applications on social robots, which can enhance the capabilities of social robots and make them more intelligent and humane (Lee et al., 2022). Social robots can use multimodal emotion recognition to sense the emotional state of the users they interact with. This includes identifying users’ facial expressions, voice emotions, text emotions, and other modal emotional signals (Laban et al., 2022). The robot can then adjust its interaction to better meet the user’s emotional needs, providing support, comfort or entertainment. In addition, social robots can use MCER to better understand users’ needs and emotional states to provide personalized suggestions and assistance. For example, emotion analysis can detect in real time the user’s anxiety, frustration, or fatigue during the interaction process. The chatbot can then adjust its language response strategy and provide more soothing tone and suggestions, thereby significantly improving the user’s trust and satisfaction. After the introduction of the emotion perception mechanism, the user’s positive feedback rate on the chatbot has been significantly improved (Islam and Bae, 2024).
8. Privacy and Security of Multimodal Data
With the widespread application of multi-modal conversational emotion recognition, privacy concerns have become increasingly critical (Yin et al., 2024). Unlike unimodal data, multi-modal data often involves sensitive personal information spanning facial expressions, voice patterns, textual content, and physiological signals. These data types may reveal not only the user’s identity but also intimate emotional states, behavioral tendencies, and even mental health conditions. Consequently, the risk of personal emotional information leakage poses a significant challenge to the secure deployment of these technologies. To address these issues, privacy protection must be considered at every stage of system design and data processing. First, data anonymization techniques should be applied to remove or obfuscate identifying information, such as name, face, or unique voice features (Ye et al., 2024a; Hanisch et al., 2025). Second, data encryption techniques (Zhang et al., 2024f; Rieyan et al., 2024) should be used to ensure protection against unauthorized access to sensitive data, whether in storage, transmission, or computation. In a distributed environment, federated learning (Zhang et al., 2024c) provides an effective framework in which model training is performed locally on the user’s device and only encrypted model updates are shared with a central server, thereby protecting the privacy of the original data. Moreover, differential privacy (Xu et al., 2025) can be introduced to inject calibrated noise into feature representations or model outputs, reducing the possibility of individual re-identification. In scenarios where multiple institutions or agents collaborate, secure multi-party computation and homomorphic encryption offer mechanisms for privacy-preserving joint model training or inference, albeit at the cost of computational efficiency (Xu et al., 2025). Finally, privacy-preserving representation learning is gaining traction, where adversarial training or disentangled learning techniques are used to suppress sensitive attributes (e.g., user identity) while preserving task-relevant emotional information (Feng et al., 2024).
@@ 9.2. Data is Heterogeneous and Noisy
MCER models need to fully eliminate heterogeneity and noise information between modalities to achieve accurate prediction or classification of emotions. Multi-modal data is naturally heterogeneous, and features of different modalities have huge differences in processing methods and representation forms. Additionally, multimodal conversation data often contains a large amount of redundant or noisy information. The emotion is typically determined by a small amount of consistent key information, such as specific words in a sentence, a particular frequency band in speech, or a distinct expression in a video. Even in some extreme cases, part of the modal information is basically unavailable under noise interference, such as ambiguous sentence expressions, noise in the speech, blocked expressions, etc. Therefore, the heterogeneity and noise of data limit the effectiveness of current multi-modal conversational emotion recognition models.
@@ 9.4. Consistent Semantic Association
Multimodal conversation emotion recognition requires the model to learn the consistent semantics across modalities to filter out noise and eliminate heterogeneity. This is essential for building an accurate multimodal emotion recognition model. However, the consistent semantic association in multimodal conversation is more complex. It is not only related to the multimodal context but also influenced by factors such as the conversation scene, the speaker’s emotional inertia, and their responses. In addition, multi-modal data are heterogeneous, each modality has differentiated representation and distribution characteristics in space, and some consistent semantic associations are hidden in the feature distribution space between modalities. Therefore, efficiently performing consistent semantic association is the primary issue that needs to be considered at the model level.
@@ 10.1. Multi-modal Conversation Data Generation
To solve the problem of ensuring strong correlation and synergy among modalities in Multi-modal Conversation Data Generation, researchers have designed advanced generative frameworks that explicitly model the cross-modal dependencies during the generation process. One representative approach leverages variational autoencoders (VAEs), which map different modalities into a shared latent semantic space, allowing the model to capture deep inter-modal relationships and generate coherent multi-modal conversational data through joint reconstruction. This shared representation ensures that generated text, audio, and visual signals are semantically aligned and contextually consistent. In addition, GAN-based frameworks introduce modality-specific discriminators alongside a joint discriminator to constrain both the individual quality and overall coherence of generated modalities. By adversarially optimizing the generator to produce modality-consistent outputs, these methods effectively enhance cross-modal correlation in generated conversations. Recently, diffusion models have shown strong potential for multi-modal data generation by modeling the complex joint distribution of multiple modalities through iterative denoising steps. Diffusion-based methods (Ho et al., 2020) can incorporate cross-modal conditional signals at each generation step, ensuring that the evolving text, audio, and visual outputs remain temporally synchronized and semantically coupled.
@@ 10.2. Multi-modal Feature Deep Fusion
To solve the above problems, on the one hand, the deformable temporal convolution can be used to allow each modality to dynamically sample the most relevant time step based on its own features. Then, the locality-aware attention is utilized to focus on the strong correlation of local areas in space. The time period misalignment and local information loss caused by heterogeneity are solved before fusion. On the other hand, to achieve cross-modal alignment, a dynamic weighted alignment mechanism can be introduced to calculate the dynamic weights of consistency and complementarity between modalities for each moment or local spatial area. Through the cross-modal attention module, the consistency score and complementarity score of each modality at the current spatiotemporal point are calculated. Then, the dynamic gating mechanism is used to adaptively adjust the contribution weight of each modal feature during fusion according to the score to avoid information redundancy or key feature loss caused by static weighting.
@@ 10.3. Unbiased Emotional Learning
To effectively solve the problem of small samples in MCER, we can start from three dimensions: enhancing samples, optimizing model structure, and adjusting training strategies, and build an integrated framework of data augmentation, prototype modeling, and category balance training. Specifically, at the data augmentation level, the emotional sentences of small sample categories are combined with context to reconstruct new samples, and the diversity of small sample categories is improved while ensuring the rationality of the context. At the prototype modeling level, a multimodal prototype center is constructed for each emotion category, and the distance between samples and prototypes of the same type is shortened by contrastive loss, and the distance from prototypes of other types is pushed away to alleviate the problem of inter-class imbalance caused by differences in sample size. At the category balance training level, the category balance focal loss or label distribution-aware margin loss is used to dynamically adjust the loss penalty items of each category. Without changing the overall training process, the model’s attention to small sample categories is improved.
@@ 10.4. Incomplete Multi-modal Conversation Emotion Recognition
To solve the common problem of missing modalities in MCER, researchers have proposed a variety of modal restoration methods in recent years, aiming to maintain the discriminative ability of the model when some modalities are unavailable. A common type of method is based on autoencoders or variational autoencoders (VAEs), which reconstruct the representation of the missing modality using the available modalities by learning the mapping relationship between different modalities. Another type of method uses generative adversarial networks (GANs) to constrain the generated features by introducing discriminators to improve the authenticity and diversity of modality completion. There are also studies that use knowledge distillation strategies to guide student models to maintain performance in modality-incomplete scenarios with the help of teacher models trained under complete modalities. Recently, the development of generative models has also provided stronger modeling capabilities for modality restoration. Diffusion models have strong distribution modeling and step-by-step optimization capabilities by modeling the reverse generation process from noise to data. They perform well in conditional generation tasks and are suitable for high-quality restoration of missing modalities. Meanwhile, flow-based models accurately model the joint distribution of modalities through reversible transformations and support conditional sampling to achieve missing modality completion.
@@ 10.5. Zero-shot Multi-modal Conversation Emotion Recognition
In recent years, large-scale pretrained models, especially cross-modal models such as CLIP (Radford et al., 2021), Flamingo (Alayrac et al., 2022), GPT-4V (Yang et al., 2023), etc., have brought a new paradigm for zero-shot multimodal emotion recognition. These models usually have strong multimodal alignment capabilities and natural language understanding capabilities, and can achieve category expansion through contrastive learning or generative modeling of text-image or text-audio. For example, CLIP builds a shared embedding space through image-text contrastive learning, which can convert emotion labels into natural language descriptions to achieve open-class emotion recognition. Audio-text models such as Whisper (Cao et al., 2012) and AudioCLIP (Guzhov et al., 2022) are also used to map speech emotion embeddings to language space to achieve zero-shot emotion transfer. In addition, the generative capabilities of multimodal large models (MLLMs) (Liu et al., 2023) can transform emotion recognition tasks into natural language generation or question-answering problems, and complete the understanding and prediction of new emotion categories under unsupervised or weak supervision. Combined with strategies such as prompt engineering, instruction tuning, or emotionally-informed prompts, the model can generalize to new emotion categories with only label descriptions.
@@ 10.6. Multi-modal Conversation Multi-label Emotion Recognition
To address the multi-label expression problem in MCER, researchers have proposed a range of targeted solutions to capture the complexity and coexistence of emotions within conversations. A common approach is to replace traditional Softmax classification with Sigmoid activation, enabling the model to independently predict the probability of each emotional label and naturally support the coexistence of multiple emotions such as sadness and anger. In addition, label dependency modeling is widely adopted, where the statistical co-occurrence patterns of emotions are explicitly captured using techniques such as label graphs or graph neural networks (GNNs), ensuring semantic consistency and reducing contradictory label outputs. To further improve robustness in real-world noisy environments, uncertainty-aware mechanisms have been integrated into multi-label frameworks, allowing the model to dynamically adjust label confidence based on the reliability of different modalities. Moreover, some studies introduce auxiliary tasks, such as emotion intensity regression or label quantity estimation, to provide richer supervision signals and enhance the model’s ability to represent complex emotional states. Recently, researchers have also explored label-specific attention mechanisms, dynamically modulating the contribution of each modality for different emotional labels, which effectively improves the multi-label recognition accuracy under modality imbalance or incomplete scenarios.
10.7. Multi-modal Emotion Recognition in Dynamic Dialogue Scenarios
Real-world conversations are inherently dynamic, with speaker roles changing depending on the context and interpersonal relationships evolving over time. However, most current multi-modal conversational emotion recognition (MCER) methods rely on static modeling frameworks that fail to capture these real-time dynamics. To address this gap, future research should focus on developing models that can adapt to such temporal and structural variability in conversation.
One promising direction is the use of time-series modeling techniques. Methods such as Temporal Convolutional Networks (TCN) (Hussain et al., 2024; Sheng et al., 2024) and Hierarchical Recurrent Neural Networks (HRNN) (Sahin and Kozat, 2024; Zhou et al., 2024) can be employed to model frame-level and utterance-level emotional fluctuations across a dialogue sequence. These models can maintain temporal dependencies and detect sudden or gradual emotional transitions by capturing long-range contextual signals. For example, by feeding sequential utterance embeddings into a TCN layer with causal convolutions, the model can learn how earlier statements influence emotional progression. Another critical aspect is the dynamic modeling of speaker interactions. Dynamic Graph Neural Networks (Dynamic GNNs) (Yang et al., 2024; Wang et al., 2024b; Zhang et al., 2024a) provide a powerful framework to represent evolving dialogue structures. Here, each node in the graph represents a speaker utterance or a speaker entity, and edges represent context-aware relationships (e.g., speaker interactions, turn-taking). Unlike static GCNs, dynamic GNNs update node and edge embeddings over time based on incoming utterances and relational changes. For instance, Temporal Graph Networks (TGNs) and EvolveGCN can learn temporal node representations by integrating recurrent modules that evolve the graph state as the dialogue progresses. These approaches allow the system to adjust for speaker role shifts and emotional influence propagation between participants. Additionally, speaker-role modeling mechanisms can be introduced by using speaker-aware encoders that condition emotion predictions on dynamic role embeddings. For example, using attention modules that incorporate speaker identity, conversational history, and position in the dialogue tree can help model role-specific emotional behaviors. Coupled with adaptive memory modules (e.g., Transformer-based memory networks), the system can retain evolving emotional cues and speaker traits across multiple turns. Finally, hybrid systems that combine reinforcement learning for emotion trend tracking and graph-based relational reasoning can adaptively adjust prediction strategies based on dialogue context evolution. This provides a more robust mechanism for handling emotional ambiguity and inter-speaker dynamics in realistic, multi-turn conversations.
10.8. Lightweight and Efficient Multimodal Conversational Emotion Recognition
Although most existing MCER models have demonstrated strong emotion recognition performance, their computational complexity and resource requirements hinder deployment in practical applications, especially on resource-constrained platforms such as mobile devices, wearable devices, and embedded systems. In real-world scenarios, real-time emotion recognition is essential for applications such as mobile assistants, edge-based healthcare monitoring, and socially interactive robots. Therefore, designing lightweight and efficient MCER models is critical for practical adoption.
To achieve real-time emotion recognition, the deployment efficiency of the model on resource-constrained devices (such as mobile terminals and embedded systems) must be solved. One of the key paths is to compress and optimize the emotion recognition model so that it can significantly reduce the computational overhead while maintaining high recognition accuracy. First, model compression can transfer the predictive ability of a large model (i.e., teacher model) trained on a high-performance platform to a smaller model (student model) with a more compact structure through methods such as knowledge distillation. During the distillation process, the student model not only learns the original labels of the training data, but also learns the ”soft labels” output by the teacher model, that is, the probability distribution of each emotion category. This method allows the small model to learn richer feature representations. The compressed model greatly reduces the number of parameters and inference delay without sacrificing too much accuracy. Secondly, pruning technology mainly evaluates the importance of certain neurons or channels in the network to the final output and prunes the less influential parts. Pruning is usually divided into two categories: structured pruning and unstructured pruning. Structured pruning can remove the entire convolution kernel, channel or layer. This method is more hardware-friendly and easy to accelerate and parallelize. Unstructured pruning performs sparse processing at the weight level. Although it has more detailed control over precision loss, it requires special sparse matrix acceleration support in actual deployment. Through multiple rounds of pruning and fine-tuning, the model complexity can be further reduced and the inference speed can be improved. Finally, the quantization method converts the original floating-point model parameters and intermediate activation values into a lower bit-width representation, such as compressing from 32-bit floating points to 8-bit integers (INT8) or even lower bit widths. Quantization can significantly reduce the storage requirements and memory bandwidth consumption of the model, while fully utilizing the integer operation acceleration capabilities of the hardware on many mobile and edge computing devices. In order to avoid the decline in recognition effect due to reduced accuracy, strategies such as quantization-aware training (QAT) or post-training quantization (PTQ) are usually used to fine-tune the weights and activation distributions to maintain the accuracy of emotion recognition.
@@ 12. Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant No. 62372478, 62472165), the YueLuShan Center Industrial Innovation (2026YCII0126), the Research Foundation of Education Bureau of Hunan Province of China (Grant No. 22B0275), the Hunan Provincial Natural Science Foundation General Project (Grant No. 2025JJ50380), and the Hunan Provincial Natural Science Foundation Youth Project (Grant No. 2025JJ60420).